OpenStudy (anonymous):

Help me find the slope of the line through this pair of points (1,-19),(-2,-7)

OpenStudy (anonymous):

Slope is rise over run, or rather, rise/run. Rise is the change in Y, and run is the change in X. \[(-7-(-19))/(-1 - 1) = 12/(-2) = -6\]

OpenStudy (raden):

-2-1 for the denominator, @kelberry